RawlsK29:

What is the value of x 9 - 2x = -8 - 4x

RawlsK29:

@Vocaloid @ThisGirlPretty

Vocaloid:

combine like terms since you have -4x on the right side of the eqn you can add 4x to both sides to get 9 - 2x + 4x = -8 - 4x + 4x the 4x's will cancel out on the right side to get 9 + 2x = -8 from there, you can subtract 9 from both sides and divide by 2 to isolate x

RawlsK29:

Oh okay, so the answer is -17/2?

Vocaloid:

good

RawlsK29:

Thanks!(: